# Replacing Softmax Similarity with a Sharpened Angular Similarity: Theory and Practice of Scaling to Billion-Context Attention

### v2

@@ 6.2 Kernel construction with the bias term
7 Causal Race Attention
@@
5:  Initialize cumulative bucket statistics:
$A_{\mathrm{cum}}^{(\ell)}\leftarrow\mathbf{0}_{R}\in\mathbb{R}^{R},\qquad B_{\mathrm{cum}}^{(\ell)}\leftarrow\mathbf{0}_{R\times d}\in\mathbb{R}^{R\times d}.$
@@
16:return $\widehat{O}=\begin{bmatrix}\widehat{O}_{1}^{\top}\\ \vdots\\ \widehat{O}_{N}^{\top}\end{bmatrix}\in\mathbb{R}^{N\times d}$.
We implemented the causal version (see Algorithm 2) efficiently using OpenMP-based parallelization rather than a naive nested-loop approach. Each hash table is processed in a separate thread with its own cumulative bucket arrays, and updates are performed incrementally in a single left-to-right scan. This avoids redundant recomputation at every step using torch.cumsum() and enables CPU-level parallel execution with negligible synchronization overhead.
Figure 6: RACE Attention pipeline from the inputs $Q,K,V\in R^{N\times d}$ to the output $O\in R^{N\times d}$: queries/keys are soft-hashed into $R$ buckets across $L$ tables and $M$ ensembles, keys/values form per-bucket summaries, and each query mixes the matched summaries to produce $O$.
Figure 7: An intuitive schematic of how RACE Attention runs with $L$ hash tables and $R$ buckets per table. Similarity between Queries and Keys is highest if they both hash to same buckets across all hash tables.

### v3

@@ 6.2 Kernel construction with the bias term
6.4 Causal Race Attention
@@
5:  Initialize cumulative bucket statistics: $A_{\mathrm{cum}}^{(\ell)}\leftarrow\mathbf{0}_{R}\in\mathbb{R}^{R}$, $B_{\mathrm{cum}}^{(\ell)}\leftarrow\mathbf{0}_{R\times d}\in\mathbb{R}^{R\times d}$.
@@
16:Return: $\widehat{O}=\begin{bmatrix}\widehat{O}_{1}^{\top}\\ \vdots\\ \widehat{O}_{N}^{\top}\end{bmatrix}\in\mathbb{R}^{N\times d}$.
We implemented the causal version (see Algorithm 2) efficiently using OpenMP/CUDA-based parallelization rather than a naive nested-loop approach. Each hash table is processed in a separate thread with its own cumulative bucket arrays, and updates are performed incrementally in a single left-to-right scan. This avoids redundant recomputation at every step using torch.cumsum() and enables CPU/GPU-level parallel execution with negligible synchronization overhead.
Figure 7: RACE Attention pipeline from the inputs $Q,K,V\in R^{N\times d}$ to the output $O\in R^{N\times d}$: queries/keys are soft-hashed into $R$ buckets across $L$ tables, keys/values form per-bucket summaries, and each query mixes the bucket summaries to produce $O$.
Figure 8: An intuitive schematic of how RACE Attention runs with $L$ hash tables and $R$ buckets per table. Similarity between Queries and Keys is highest if they both assign a higher mass to same buckets across all hash tables.